OpenStudy (anonymous):

During the month of February, Fabulous Feet Shoe Mart sold 50 pairs of red loafers. After an ad campaign to boost sales, they sold 60 pairs in March. Find the percent of increase in sales.

OpenStudy (mokeira):

Increase= 60-50 Increase= 10 Percentage increase= \[\frac{ 10 }{ 100 }\times100\] 10% increase